# Kitashirakawa Oiwakecho, Sakyo-ku, Kyoto 606-8502, Japan

### v1

@@ 5 Causality on the holographic screen
In this section, we present an apparent failure of the connected wedge theorem in dS3 spacetime. As the theorem has a proof based on quantum information, which should apply in any holographic setup, this poses a puzzle. We resolve it by revisiting causality on the holographic screen. This resolution leads to three important consequences: 1) An insight for the dS connected wedge theorem, proved in Section 6 for three-dimensional asymptotically dS spacetime. 2) Bulk local excitations in the interior of an observer’s holographic screen, emanating from the screen, are not described by local operators on the screen. 3) A local excitation on $\mathcal{S}$ should be mapped to a local operator on $\mathcal{I}^{\pm}$, hinting at a relation between static patch holography and the dS/CFT correspondence.
@@ 5.1 An apparent violation of the connected wedge theorem
To this end, one needs to identify the appropriate causality on the screen $\mathcal{S}$. Let us consider some geometrically reasonable choices of causality, which do not rely on the details of the microscopic theory.777or a proposal of a possible microscopic resolution such as wormholes, see Omiya:2021olc . One possibility would be the causality based on the induced metric on $\mathcal{S}$. In this case, the screen causality is completely blind to the holographic bulk. In the current limiting case, the discrepancy between the bulk causality and boundary causality is the highest, as $\mathcal{S}$ is located on the cosmological horizon, i.e. the induced metric is
@@
Thus, under this causality, only light can propagate at a fixed angle $\varphi$. This suggests that the decision region is empty as the spatial location of each input and output point differs. [^8] There is a subtlety at $\sigma=0$ as the cosmological horizon bifurcates. However, this subtlety can be avoided without significantly modifying the metric by replacing $\mathcal{S}$ with the stretched horizon. Thus, the conclusion should be the same. Note that the decision region becomes nonempty only when the spatial location of one of the input points and both output points coincide. Even if this happens, the decision region is pointlike so it leads to a significant violation of the connected wedge theorem.
@@ 5.2 Induced causality from dS boundary $\mathcal{I}^{\pm}$
What needs to be modified in accordance with the connected wedge theorem? Since we start from the semiclassical limit of the holographic duality, the existence of the bulk scattering from the bulk causality should be taken for granted. On the other hand, there is room to change the causality and decision regions fixed from it on the screen where the dual theory lives. We implicitly assumed through the definition of the decision regions that a bulk signal emanating from or reaching a point on the screen is given by an excitation of a local operator on the screen. This is the case in AdS/CFT, where a bulk excitation can be created by a local operator on the asymptotic boundary Nozaki:2013wia ; Terashima:2023mcr . It was noted by one of the authors in Mori:2023swn that this is not the case in AdS holography with braneworld or cutoff surfaces. They suggest that a correct boundary dual of a localized wave packet in the bulk is given by a local excitation on a fictitious boundary so that it is effectively smeared and nonlocal on the brane/cutoff surface. [^9] Alternatively, this is interpreted as preparing the operator by some finite Lorentzian time evolution. This hints that we need to define another notion of causality on the non-asymptotic boundary, namely, the induced causality.
In this work, we follow the strategy of Mori:2023swn and consider ‘fictitious’ local perturbations at the conformal boundaries of asymptotically de Sitter spacetime. We quote the word ‘fictitious’ because in dS the screen is not a boundary, where spacetime terminates, so the conformal boundary is not fictitious contrary to the previous work in AdS. This will be first motivated by the analogy with the case studied in Mori:2023swn , reviewed in Section 3, and related to geometric properties of causal horizons, which will be made precise in Section 6.2. The second motivation is the dS/CFT correspondence Strominger:2001pn , which would provide a great physical interpretation of the fictitious points providing the induced lightcones of perturbations on the holographic screen.
@@
Definition 5.1 (‘Fictitious’/tilted point).
@@
The tilted point associated with a point on the screen leads to the definition of an induced lightcone:
@@
This gives the formula for the induced lightcone on a screen at a constant $r$ in the static patch in the conformal coordinates $(\sigma,\theta=\pi-\arcsin(r\cos\sigma),\varphi)$ as a function of the location of the tilted point $(\sigma_{i}=k(\pi/2-\epsilon_{dS}),\theta_{i},\varphi_{i})$ at $\mathcal{I}^{\pm}$. [^10] We define the domain of the inverse sine function as $\arcsin x\in[-\pi/2,\pi/2]$.
@@ 6.1 Resolution of the apparent contradiction
A direct scattering on $\mathcal{S}$ is not possible from the induced screen causality, as $\mathcal{R}_{1}\cap\mathcal{R}_{2}=\varnothing$. Now that $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ are specified, their entanglement entropy can be computed. As reviewed in Appendix A.1, the entanglement entropy of a spatial subsystem of the screen $\mathcal{S}$ has three contributions. The entropy is given by the sum of the areas of the homologous minimal extremal surfaces 1) in the interior of $\mathcal{S}$, 2) in the interior of a complementary screen associated with an antipodal observer, and 3) in the exterior region bounded by the union of these two screens. These three regions are illustrated in gray in Figure 10. We denote $D$ the causal diamond inside $\mathcal{S}$ in which we extremize the area.
@@
$\gamma_{e}^{dis}(\mathcal{V}_{1}\cup\mathcal{V}_{2},D)=\gamma_{e}(\mathcal{V}_{1},D)\cup\gamma_{e}(\mathcal{V}_{2},D),$ (40)
where $\gamma_{e}(\mathcal{V}_{i},D)$ is the geodesic anchored to $\mathcal{V}_{i}$ in the region inside $\mathcal{S}$, whose length is $\min(\Delta\varphi,2\pi-\Delta\varphi)$ where $\Delta\varphi$ is the angle of the arc. [^12] See appendix B.1 of Franken:2023pni for a detailed computation. See Figure 14 for the extremal surface and the homology region in the interior in each case.
@@
$\gamma_{e}^{con}(\mathcal{V}_{1}\cup\mathcal{V}_{2},D)=\gamma_{e}^{dis}((\mathcal{V}_{1}\cup\mathcal{V}_{2})^{c},D),$ (41)
@@
${\rm area}(\gamma_{e}^{dis}(\mathcal{V}_{1}\cup\mathcal{V}_{2},D_{L}))-{\rm area}(\gamma_{e}^{con}(\mathcal{V}_{1}\cup\mathcal{V}_{2},D_{L}))=0.$ (42)
In other words, the scattering we considered corresponds to the transition case where the disconnected and connected entanglement wedges are equivalent. Even though we considered input and output points on a screen located on the cosmological horizon, this result would have been identical for any set of input/output points on an arbitrary stretched horizon at fixed radius $r$ such that $\tilde{c}_{i}$ and $\tilde{r}_{i}$ of equation (37) are their associated set of tilted points on $\mathcal{I}^{\pm}$. Indeed, for these tilted points, equation (31) becomes linear in $r$, such that the size of $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ is independent of $r$. This is pictured in Figure 6. Hence, the fact that we found an exact match between the lengths of the connected and disconnected geodesics for these induced points is not specific to a screen on the horizon, but a general property of these ‘fictitious’ points that are associated with a $J_{12\rightarrow 12}$ reducing to a point. Too strong. Just suggests. This is consistent with the connected wedge theorem, and the typical behavior of pointlike holographic scatterings where the lengths of the connected and disconnected geodesics are equivalent May:2019odp ; May:2019yxi ; May:2021nrl ; May:2022clu . This example thus provides good evidence that the induced lightcones defined in Section 5.2 are the right objects to consider when studying causality on the holographic screen.
@@ 6.2 Geometric proof
  • In the bulk region inside $\mathcal{S}$, minimal extremal surfaces are maximin surfaces and conversely. We develop this statement in Theorem A.5 of Appendix A.3.
@@
Let $\mathcal{V}_{i}$ be the achronal codimension-two surface on $\mathcal{S}$ such that $\hat{D}(\mathcal{V}_{i})=\mathcal{R}_{i}$, as in Definition 6.1. $\Sigma|_{\mathcal{S}}$ denotes an achronal slice of $\mathcal{S}$ and $\Sigma$ denotes an achronal slice containing the observer of interest in the bulk such that $\partial\Sigma=\Sigma|_{\mathcal{S}}$. $D=D(\Sigma)$ denotes the causal diamond of slices in the static patch that is bounded by $\Sigma|_{\mathcal{S}}$. Note that the definition of $\Sigma$ here is either $\Sigma_{L}$ or $\Sigma_{R}$, which is different from the convention used in Appendix A. The exact location of the complementary screen $\mathcal{S}_{R}$ is not important here. Let us denote $\gamma_{i}=\gamma_{e}(\mathcal{V}_{i},D)$ with $i=1,2$ and $\gamma_{1\cup 2}=\gamma_{e}(\mathcal{V}_{1}\cup\mathcal{V}_{2},D)$ for simplification. To show Theorem 6.1, we follow the strategy of May:2019odp . We need to prove that
@@
where $\mathcal{C}_{1},\mathcal{C}_{2}\in\Sigma$ are the homology regions associated with $\gamma_{1}$ and $\gamma_{2}$ on $\Sigma$ [^17] The homology region is a spacelike slice bounded by the extremal surface and its associated subsystem of the screen. Its causal diamond gives the corresponding entanglement wedge. See Definition A.1. $J_{12\rightarrow 12}$ is assumed to be non-empty, such that the right-hand side of equation (48) is also non-empty. This region being not empty implies that a subset of $J^{+}(\mathcal{C}_{1}\cup\mathcal{C}_{2})$ exists in the past of $\tilde{r}_{1}$ and $\tilde{r}_{2}$. Because $J^{+}(\mathcal{C}_{1}\cup\mathcal{C}_{2})=J^{+}(\mathcal{C}_{1})\cap J^{+}(\mathcal{C}_{2})$, $J^{+}(\mathcal{C}_{1})$ and $J^{+}(\mathcal{C}_{2})$ must meet in the past of $\tilde{r}_{1}$ and $\tilde{r}_{2}$. The ridge is therefore non-empty.
@@
The slope is a subsystem of two causal horizons associated with $\tilde{r}_{1}$ and $\tilde{r}_{2}$ on $\mathcal{I}^{+}$, which is in the future of $\Sigma$ and in the past of the lightsheets emanating from $\gamma_{1}$ and $\gamma_{2}$. We now construct the null membrane as
@@
We have constructed on every slice $\Sigma$ a codimension-two surface homologous to $\mathcal{V}_{1}\cup\mathcal{V}_{2}$ with a smaller area than the extremal area surface $\gamma_{1}\cup\gamma_{2}$ with a disconnected entanglement wedge, implying that $\gamma_{1}\cup\gamma_{2}$ is not a true maximin surface, since it is not minimal on any slice $\Sigma$. Minimal extremal surfaces in the interior of $D$ are maximin surfaces (Theorem A.5) so $\gamma_{1}\cup\gamma_{2}$ cannot be the smallest extremal surface, concluding the proof. ∎
@@ 7 Discussion
In this paper, we did not discuss the detailed nature of a local excitation at null infinity or ‘smeared’ excitation on the screen. In fact, knowing what types of excitations are allowed is very important to support our proof of the connected wedge theorem Mori:2023swn . Nevertheless, at this point the precise holographic dual of de Sitter, in particular within static patch holography, is unclear. The DSSYK model Susskind:2021esx ; Lin:2022nss ; Rahman:2022jsf ; Goel:2023svz ; Narovlansky:2023lfz ; Verlinde:2024znh ; Verlinde:2024zrh ; Blommaert:2023opb ; Blommaert:2023wad ; Rahman:2024iiu is one candidate, however, the precise location of the screen is still debatable. Additionally, it is dimensionally reduced so the dual of dS3 is further unclear. In the proposed dS/DSSYK duality, Susskind has argued that only a small number of special collective excitations of the fermions in the DSSYK model would propagate deep in the bulk Susskind:2022bia . It is interesting to see what type of excitations in the theory realize the holographic scattering if there are, and see if the model shows signs of induced causality. Furthermore, as we mentioned in the introduction, a further investigation of the connected wedge theorem in a concrete theory could help us to identify the true UV boundary, from which holographic dS emerges.
Another possible direction is to construct a finite-dimensional toy model. Since quantum tasks in nonlocal quantum computations are better understood in finite dimensions, often by using circuit diagrams, we may be able to discuss a general feature of holographic quantum tasks by modeling static patch holography by a (random) quantum error correcting code as in AdS/CFT . This also naturally provides a holographic dictionary that maps the effective picture, where a local unitary via direct scattering takes place, to the fundamental picture, where the nonlocal quantum computation takes place, and vice versa Akers:2022qdl ; Akers:2021fut ; Akers:2022qdl .
@@
An implicit assumption made in the connected wedge theorem is that the evolution on the holographic screen(s) is local. Otherwise, the time evolution manifestly entangles two decision regions so it cannot be viewed as a nonlocal quantum computation. [^20] We thank Beni Yoshida for pointing this out. The Hamiltonian generating time evolution in the two static patches is $H=H_{R}-H_{L}$, where $H_{L,R}$ are the Hamiltonian for each static patch. On the other hand, the screen dual of the evolution in the Milne patch toward $\mathcal{I}^{+}$ would be $H=H_{R}+H_{L}$. Similar to the eternal AdS black hole, this modular evolution may couple the left and the right screen. Thus, it is essential to understand the locality of the time evolution on the screens to justify the connected wedge theorem in the exterior region.
@@
In the de Sitter version of the connected wedge theorem proved in this paper, we related input and output points to points at past and future null infinities. This maps a bulk scattering from the static patch holographic screens to a bulk scattering from conformal boundaries. It would be interesting if the dS/CFT correspondence could provide a precise framework to describe this type of scattering. In particular, can we prove a connected wedge theorem in dS/CFT? Doing so would require precise notions of causality on the Euclidean CFT, which is problematic as there is no notion of time evolution. Our construction may serve as an alternative definition of ‘time’ in dS/CFT as it provides a mapping between points on the static patch, which follows the real time evolution, to the points on the conformal boundary. Additionally, studying connected wedges in dS/CFT would require a precise prescription to compute entanglement wedges of CFT subregions. This has been studied in Doi:2022iyj ; Narayan:2022afv ; Narayan:2023zen ; Doi:2023zaf ; Das:2023yyl ; Doi:2024nty . Succeeding in this challenge would open a path to a better understanding of time, entanglement, and causality in dS/CFT.

### v3

@@ 5 Causality on the holographic screen
In this section, we present an apparent failure of the connected wedge theorem in dS3 spacetime. As the theorem has a proof based on quantum information, which should apply in any holographic setup, this poses a puzzle. (Note that the quantum information argument implicitly requires unitarity. See a remark after this paragraph.) We resolve it by revisiting causality on the holographic screen. This resolution leads to three important consequences: 1) An insight for the dS connected wedge theorem, presented in Section 6 for three-dimensional asymptotically dS spacetime. 2) Bulk local excitations in the interior of an observer’s holographic screen, emanating from the screen, are not described by local operators on the screen. 3) A local excitation on $\mathcal{S}$ should be mapped to a local operator on $\mathcal{I}^{\pm}$, hinting at a relation between static patch holography and the dS/CFT correspondence.
Remark 2.
In Section 3.1, we implicitly assume unitarity on the screen/boundary in the quantum information argument for the connected wedge theorem. One may think this is not the case for a static patch of dS as it is a part of the global dS. We argue however that this would be the case. Recall that there is a timelike Killing vector in a static patch of pure dS. This implies that there is a Hamiltonian that generates a time translation so the state in the static patch is expected to evolve unitarily, according to the Liouville-von Neumann equation. This does not contradict with the static patch being a part of the global spacetime since the state can be mixed; in fact, it should be an (almost) maximally mixed state carrying dS entropy. We should stress that, however, a static patch may not have a timelike Killing vector in asymptotically dS, as mentioned in footnote 2. Nevertheless, our gravitational proof that we present later in Section 6.2 works even in these cases by defining the static patch as the causal region (timelike envelope) of the observer’s worldline.
@@ 5.1 An apparent violation of the connected wedge theorem
To this end, one needs to identify the appropriate causality on the screen $\mathcal{S}$. One possibility would be the causality based on the induced metric on $\mathcal{S}$. In this case, the screen causality is completely blind to the holographic bulk. In the current limiting case, the discrepancy between the bulk causality and boundary causality is the highest, as $\mathcal{S}$ is located on the cosmological horizon, i.e. the induced metric is
@@
Thus, under this causality, only light can propagate at a fixed angle $\varphi$. This suggests that the decision region is empty as the spatial location of each input and output point differs. [^14] There is a subtlety at $\sigma=0$ as the cosmological horizon bifurcates. However, this subtlety can be avoided without significantly modifying the metric by replacing $\mathcal{S}$ with the stretched horizon. Thus, the conclusion should be the same. Note that the decision region becomes nonempty only when the spatial location of one of the input points and both output points coincide. Even if this happens, the decision region is pointlike so it leads to a significant violation of the connected wedge theorem. Indeed, equation (20) leads to a vanishing mutual information since the area of $\gamma_{e}(A;D)$ goes to 00 as $A$ becomes pointlike. The disconnected extremal surfaces $\gamma_{e}(A;D)\cup\gamma_{e}(B;D)$ therefore always dominate when $A$ and $B$ are pointlike. See Appendix B for more details on the holographic entanglement entropy prescription.
@@ 5.2 Induced causality from dS boundary $\mathcal{I}^{\pm}$
What needs to be modified for the connected wedge theorem to hold? Since we start from the semiclassical limit of the holographic duality, the existence of the bulk scattering from the bulk causality should be taken for granted. On the other hand, there is room to change the causality and decision regions fixed from it on the screen where the dual theory lives. We implicitly assumed through the definition of the decision regions that a bulk signal emanating from or reaching a point on the screen is given by an excitation of a local operator on the screen. This is the case in AdS/CFT, where a bulk excitation can be created by a local operator on the asymptotic boundary Nozaki:2013wia ; Terashima:2023mcr . It was noted by one of the authors in Mori:2023swn that this is not the case in AdS holography with braneworld or cutoff surfaces. They suggest that a correct boundary dual of a localized wave packet in the bulk is given by a local excitation on a fictitious boundary so that it is effectively smeared and nonlocal on the brane/cutoff surface. [^15] Quantum mechanically, the fictitious evolution is interpreted as a state preparation on the hypersurface by a finite Lorentzian time evolution. Alternatively, a fine-tuned causal structure from microscopic aspects of the dual theory may resolve the puzzle Omiya:2021olc . In this paper, we seek a resolution that only relies on the geometry, without such a fine tuning of the theory. This hints that we need to define another notion of causality on the non-asymptotic boundary, namely, the induced causality.
In this work, we follow the strategy of Mori:2023swn and consider ‘fictitious’ local perturbations at the conformal boundaries of asymptotically de Sitter spacetime. We quote the word ‘fictitious’ because in dS the screen is not a boundary, where spacetime terminates, so the conformal boundary is not fictitious contrary to the previous work in AdS. This will be first motivated by the analogy with the case studied in Mori:2023swn , reviewed in Section 3, and related to geometric properties of causal horizons, which will be made precise in Section 6.2. The second motivation is the dS/CFT correspondence Strominger:2001pn , which would provide a physical interpretation of the fictitious points providing the induced lightcones of perturbations on the holographic screen.
@@
Definition 5.1 (‘Fictitious’/tilded point).
@@
The tilded point associated with a point on the screen leads to the definition of an induced lightcone:
@@
This gives the formula for the induced lightcone on a screen at a constant $r$ in the static patch in the conformal coordinates $(\sigma,\theta=\pi-\arcsin(r\cos\sigma),\varphi)$ as a function of the location of the tilded point $(\sigma_{i}=k(\pi/2-\epsilon_{dS}),\theta_{i},\varphi_{i})$ at $\mathcal{I}^{\pm}$. [^16] We define the domain of the inverse sine function as $\arcsin x\in[-\pi/2,\pi/2]$.
@@ 6.1 Resolution of the apparent contradiction
A direct scattering on $\mathcal{S}$ is not possible from the induced screen causality, as $\mathcal{R}_{1}\cap\mathcal{R}_{2}=\varnothing$. Now that $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ are specified, their entanglement entropy can be computed. As reviewed in Appendix B.1, the entanglement entropy of a spatial subsystem of the screen $\mathcal{S}$ has three contributions. The entropy is given by the sum of the areas of the homologous minimal extremal surfaces 1) in the interior of $\mathcal{S}$, 2) in the interior of a complementary screen associated with an antipodal observer, and 3) in the exterior region bounded by the union of these two screens. These three regions are illustrated in gray in Figure 12. We denote $D$ the causal diamond inside $\mathcal{S}$ in which we extremize the area.
@@
$\gamma_{e}^{dis}(\mathcal{V}_{1}\cup\mathcal{V}_{2};D)=\gamma_{e}(\mathcal{V}_{1};D)\cup\gamma_{e}(\mathcal{V}_{2};D),$ (42)
where $\gamma_{e}(\mathcal{V}_{i};D)$ is the geodesic anchored to $\mathcal{V}_{i}$ in the region inside $\mathcal{S}$, whose length is $\min(\Delta\varphi,2\pi-\Delta\varphi)$ where $\Delta\varphi$ is the angle of the arc. [^18] See appendix B.1 of Franken:2023pni for a detailed computation. See Figure 20 for the extremal surface and the homology region in the interior in each case.
@@
$\gamma_{e}^{con}(\mathcal{V}_{1}\cup\mathcal{V}_{2};D)=\gamma_{e}^{dis}((\mathcal{V}_{1}\cup\mathcal{V}_{2})^{c};D),$ (43)
@@
${\rm area}(\gamma_{e}^{dis}(\mathcal{V}_{1}\cup\mathcal{V}_{2};D_{L}))-{\rm area}(\gamma_{e}^{con}(\mathcal{V}_{1}\cup\mathcal{V}_{2};D_{L}))=0.$ (44)
In other words, the scattering we considered corresponds to the transition case where the disconnected and connected entanglement wedges are equivalent. Even though we considered input and output points on a screen located on the cosmological horizon, this result would have been identical for any set of input/output points on an arbitrary stretched horizon at fixed radius $r$ such that $\tilde{c}_{i}$ and $\tilde{r}_{i}$ of equation (39) are their associated set of tilded points on $\mathcal{I}^{\pm}$. Indeed, for these tilded points, equation (33) becomes linear in $r$, such that the size of $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$ is independent of $r$. This is pictured in Figure 8. Hence, the fact that we found an exact match between the lengths of the connected and disconnected geodesics for these induced points is not specific to a screen on the horizon, but a general property of these ‘fictitious’ points that are associated with a $J_{12\rightarrow 12}$ reducing to a point. This is consistent with the connected wedge theorem, and the typical behavior of pointlike holographic scatterings where the lengths of the connected and disconnected geodesics are equivalent May:2019odp ; May:2019yxi ; May:2021nrl ; May:2022clu . This example thus provides evidence that the induced lightcones defined in Section 5.2 are the right objects to consider when studying causality on the holographic screen.
@@ 6.2 Geometric proof
  • In the bulk region inside $\mathcal{S}$, minimal extremal surfaces are maximin surfaces and conversely. We develop this statement in Theorem B.5 of Appendix B.3.
@@
Let $\mathcal{V}_{i}$ be the achronal codimension-two surface on $\mathcal{S}$ such that $\hat{D}(\mathcal{V}_{i})=\mathcal{R}_{i}$, as in Definition 6.1. $\Sigma|_{\mathcal{S}}$ denotes an achronal slice of $\mathcal{S}$ and $\Sigma$ denotes an achronal slice containing the observer of interest in the bulk such that $\partial\Sigma=\Sigma|_{\mathcal{S}}$. $D=D(\Sigma)$ denotes the causal diamond of slices in the static patch that is bounded by $\Sigma|_{\mathcal{S}}$. Note that the definition of $\Sigma$ here is either $\Sigma_{L}$ or $\Sigma_{R}$, which is different from the convention used in Appendix B. The exact location of the complementary screen $\mathcal{S}_{R}$ is not important here. Additionally, the exterior contribution to the entropy is irrelevant here as the exterior entanglement wedge is always disconnected. The contribution from the complementary patch is absent because the decision regions do not lie on the complementary screen. For a more detailed discussion, see Section 6.1. Let us denote $\gamma_{i}=\gamma_{e}(\mathcal{V}_{i};D)$ with $i=1,2$ and $\gamma_{1\cup 2}=\gamma_{e}(\mathcal{V}_{1}\cup\mathcal{V}_{2};D)$ for simplification. To show Theorem 6.1, we follow the strategy of May:2019odp . We need to prove that
@@
where $\mathcal{C}_{1},\mathcal{C}_{2}\in\Sigma$ are the homology regions associated with $\gamma_{1}$ and $\gamma_{2}$ on $\Sigma$ [^23] The homology region is a spacelike slice bounded by the extremal surface and its associated subsystem of the screen. Its causal diamond gives the corresponding entanglement wedge. See Definition B.1. $J_{12\rightarrow 12}$ is assumed to be non-empty, such that the right-hand side of equation (50) is also non-empty. This region being not empty implies that a subset of $J^{+}(\mathcal{C}_{1}\cup\mathcal{C}_{2})$ exists in the past of $\tilde{r}_{1}$ and $\tilde{r}_{2}$. Because $J^{+}(\mathcal{C}_{1}\cup\mathcal{C}_{2})=J^{+}(\mathcal{C}_{1})\cap J^{+}(\mathcal{C}_{2})$, $J^{+}(\mathcal{C}_{1})$ and $J^{+}(\mathcal{C}_{2})$ must meet in the past of $\tilde{r}_{1}$ and $\tilde{r}_{2}$. The ridge is therefore non-empty.
@@
The slope is a subsystem of two causal horizons associated with $\tilde{r}_{1}$ and $\tilde{r}_{2}$ on $\mathcal{I}^{+}$, which is in the future of $\Sigma$ and the past of the lightsheets emanating from $\gamma_{1}$ and $\gamma_{2}$. We now construct the null membrane as
@@
We have constructed on every slice $\Sigma$ a codimension-two surface homologous to $\mathcal{V}_{1}\cup\mathcal{V}_{2}$ with a smaller area than the extremal area surface $\gamma_{1}\cup\gamma_{2}$ with a disconnected entanglement wedge, implying that $\gamma_{1}\cup\gamma_{2}$ is not a true maximin surface, since it is not minimal on any slice $\Sigma$. Minimal extremal surfaces in the interior of $D$ are maximin surfaces (Theorem B.5) so $\gamma_{1}\cup\gamma_{2}$ cannot be the smallest extremal surface, concluding the proof. ∎
@@ 7 Discussion
In this paper, we did not discuss the detailed nature of a local excitation at null infinity or ‘smeared’ excitation on the screen. In fact, knowing what types of excitations are allowed is very important to support our proof of the connected wedge theorem Mori:2023swn . Nevertheless, at this point the precise holographic dual of de Sitter, in particular within static patch holography, is unclear. The DSSYK model Susskind:2021esx ; Lin:2022nss ; Rahman:2022jsf ; Goel:2023svz ; Narovlansky:2023lfz ; Verlinde:2024znh ; Verlinde:2024zrh ; Blommaert:2023opb ; Blommaert:2023wad ; Rahman:2024iiu is one candidate, however, the precise location of the screen is still debatable. Additionally, it is dimensionally reduced so the dual of dS3 is further unclear. In the proposed dS/DSSYK duality, Susskind has argued that only a small number of special collective excitations of the fermions in the DSSYK model would propagate deep in the bulk Susskind:2022bia . It is interesting to see what type of excitations in the theory realize holographic scattering if there are, and see if the model shows signs of induced causality. We can already point out that the induced lightcones constructed in this paper showcase a difference between perturbations propagating into the bulk and perturbations that are confined close to the stretched horizon. Indeed, the induced lightcone of a bulk propagating perturbation is apparently superluminal while it does not violate the microcausality Mori:2023swn . This means such an excitation is nonlocal and obeys a nonlocal evolution from the screen point of view. On the other hand, the lightcone of a perturbation confined on the screen is a lightcone based on the induced metric, and it does not exhibit superluminality. A further investigation of these perturbations and their duals in a concrete theory could help us to identify the true UV boundary, from which holographic dS emerges.
Another possible direction is to construct a finite-dimensional toy model. Since quantum tasks in nonlocal quantum computations are better understood in finite dimensions, often by using circuit diagrams, we may be able to discuss a general feature of holographic quantum tasks by modeling static patch holography by a (random) quantum error correcting code as in AdS/CFT . This also naturally provides a holographic dictionary between the effective picture, where a local unitary via a direct scattering takes place, and the fundamental picture, where the nonlocal quantum computation may take place Akers:2022qdl ; Akers:2021fut .
@@
An implicit assumption made in the connected wedge theorem is that the evolution on the holographic screen(s) is local. Otherwise, the time evolution manifestly entangles two decision regions so it cannot be viewed as a nonlocal quantum computation. [^26] We thank Beni Yoshida for pointing this out. The Hamiltonian generating time evolution in the two static patches is $H=H_{R}-H_{L}$, where $H_{L,R}$ are the Hamiltonian for each static patch. On the other hand, the screen dual of the evolution in the Milne patch toward $\mathcal{I}^{+}$ may require coupling the left and the right boundaries similar to the eternal AdS black hole. Thus, it is essential to understand the locality of the time evolution on the screens to justify the connected wedge theorem in the exterior region.
@@
In the de Sitter version of the connected wedge theorem argued in this paper, we related input and output points to points at past and future null infinities. This maps a bulk scattering from the static patch holographic screens to a bulk scattering from conformal boundaries. It would be interesting if the dS/CFT correspondence could provide a precise framework to describe this type of scattering. In particular, can we prove a connected wedge theorem in dS/CFT? Doing so would require precise notions of causality on the Euclidean CFT, which is problematic as there is no notion of time evolution. Our construction may serve as an alternative definition of ‘time’ in dS/CFT as it provides a mapping between points on the static patch, which follows the real time evolution, to the points on the conformal boundary. See Figure 11 for its picture. Additionally, studying connected wedges in dS/CFT would require a precise prescription to compute the entanglement entropy of CFT subregions. This has been studied in Doi:2022iyj ; Narayan:2015vda ; Narayan:2020nsc ; Narayan:2022afv ; Narayan:2023zen ; Doi:2023zaf ; Das:2023yyl ; Doi:2024nty . Succeeding in this challenge would open a path to a better understanding of time, entanglement, and causality in dS/CFT.
@@
Note added – The mapping of ‘time’ between an observer worldline in static patch and null infinity is argued in Parikh:2002py . [^27] We thank Zixia Wei for introducing this reference and a relevant discussion. The authors focus on flat slicing by starting from dS/$\mathbb{Z}_{2}$, where the quotient is given by the antipodal identification. The ‘time’ in null infinity is obtained by analytically continuing the timelike Killing vector to the Milne patch. As also mentioned by Strominger Strominger:2001pn , this flow corresponds to the scale transformation on null infinity. This gives an interpretation of the state evolution in light of the radial quantization. The setup in Parikh:2002py is different from ours, as dS is a quotient or global and the mapping is to the observer’s worldline or a holographic screen. When we focus on the ingoing or outgoing mode of scattering, the mapping only requires flat slicing, thus for each mode, there is no essential difference. However, to argue the decision regions relevant to the connected wedge, we need to combine the future and past induced light cones. This requires us to consider global dS, not a quotient one. Nevertheless, the mapping itself is the same and our discussion on the connected wedge theorem gives another evidence for the mapping proposal.